I would really appreciate it if I could get some help on this question, thanks.
The path of a falling object is given by the function s=-16t^2+vo+s0 where vo represents the initial velocity in ft/sec and so represents the initial height in feet. Also, s represents the height in feet of the object at any time,t,which is measured in seconds.
If a rock is thrown upward with the initial velocity of 40ft per second from the top of a 30ft building.
s= -16t^2+vot+so I figured out that much, then it asks after how many seconds will the graph reach maximum height? I just need help setting up the equation to solve for height....The question comes from my online algebra class at aiu. HELP PLEASE. thanks, alaska

Answer by scott8148(6628)

the maximum (or minimum) of a parabola is the vertex, which lies on the axis of symmetry

for a parabola of the form ax^2+bx+c, the equation for the axis of symmetry is x=-b/(2a)

so the time at maximum will be t=-40/(2(-16)) __ t=-40/(-32) __ t=1.25